OpenStudy (anonymous):

Find the measure of angle x. Round your answer to the nearest hundredth. (please type the numerical answer only)

OpenStudy (anonymous):

Sin cos tan- but i need help understanding

OpenStudy (anonymous):

@Hero @dan815

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

first you need the hypotenuse which you find via pythagora

OpenStudy (anonymous):

\[8^2+15^2=h^2\] solve for \(h\)

OpenStudy (anonymous):

ok hold on lemme get it

OpenStudy (anonymous):

then you can take any trig ratio you like

OpenStudy (anonymous):

17? @satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so that is the answer? bcuz it says round to the nearest hundredth.

OpenStudy (anonymous):



OpenStudy (anonymous):

that is the answer to the question "what is the hypotenuse" it is not the answer to sine, cosine or tangent

OpenStudy (anonymous):

oh ok xp

OpenStudy (anonymous):

\[\sin(x)=\frac{opp}{hyp}=\frac{8}{17}\] look up the other trig ratios so know can memorize them

OpenStudy (anonymous):

soh cah toah? or sin cos tan?

OpenStudy (anonymous):

@satellite73